# Supervised Learning-enhanced Multi-Group Actor Critic for Live Stream Allocation in Feed

### v2

@@ Abstract.
Reinforcement Learning (RL) has been widely applied in recommendation systems to capture long-term user engagement, thus improving dwelling time and improving user retention. In the context of a short video & live stream mixed recommendation scenario, the live stream recommendation system (RS) decides whether to inject at most one live stream into the video feed for each user request. To maximize long-term user engagement, it is crucial to determine an optimal live stream injection policy for accurate live stream allocation. However, traditional RL algorithms often face divergence and instability problems, and these issues may cause too many live stream allocation, which interrupts user’s short video interest and leads to a decrease in the user’s app usage duration. To address these challenges, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm (SL-MGAC). Specifically, we introduce a supervised learning-enhanced actor-critic framework that incorporates variance reduction techniques, where multi-task reward learning helps restrict bootstrapping error accumulation during critic learning. Additionally, we design a multi-group state decomposition module for both actor and critic networks to reduce prediction variance and improve model stability. We also propose a novel reward function to prevent overly greedy live stream allocation. Empirically, we evaluate the SL-MGAC algorithm using offline policy evaluation (OPE) and online A/B testing. Experimental results demonstrate that the proposed method not only outperforms baseline methods but also exhibits enhanced stability in online recommendation scenarios.
@@ 1. Introduction
Deep reinforcement learning has shown great potential in learning policies to maximize long-term rewards in various research domains, such as computer vision, robotics, natural language processing, gaming, and recommendation systems (Afsar et al., 2022; Li, 2017). In the context of recommendation systems, reinforcement learning is applied to optimize long-term user engagement (Zou et al., 2019) and improve user retention (Cai et al., 2023). Numerous RL applications have been proposed for real-world recommendation systems, including slate recommendation (Ie et al., 2019; Deffayet et al., 2023; Liu et al., 2023), personalized search ranking (Miao et al., 2021), and advertisement allocation (Liao et al., 2022).
We consider a challenging sequential decision making scenario in a short video & live stream blended recommendation system, as shown in Fig. 1. The system consists of three components: a live streaming recommendation system, a short video recommendation system, and a blending server. For each user request with timestamp $t$, the live stream recommendation system decides whether to inject the recommended live stream into the video feed, while the short video recommendation system suggests $B$ (with $B<10$) videos. The blending server then mixes and rearranges the single live stream with the $B$ videos to form the final recommendation list. Our objective is to find a personalized optimal live stream injection policy that not only maximizes user long-term engagement with live streams, but also meets the constraint of not causing a negative impact on the app usage duration, which can be modeled as an infinite request-level Constrained Markov Decision Process (CMDP) (Altman, 2021).
However, the intrinsic issues of divergence and instability associated with traditional reinforcement learning algorithms (RL) (François-Lavet et al., 2015; Fujimoto et al., 2018; Kumar et al., 2019; Dasagi et al., 2019) are significantly exacerbated in this challenging scenario. A primary possible reason for this is the drastic fluctuations in both the live stream supply scale and user interaction behaviors over time, as shown in Fig. 7 in Appendix B. RL models often struggle to learn effective policies from data that exhibit such high variance across both time scales and user scales. In practice, we observe that RL models frequently encounter issues of policy deterioration or model collapse (Dohare et al., 2023).
@@
The main contributions of this paper are as follows:
  • We introduce SL-MGAC, a novel RL model designed to reduce prediction variance while enhancing the stability of both offline training and online inference.
@@ 2. Problem Formulation
As shown in Fig. 1, in the short video & live stream mixed recommendation system, the live stream recommendation system is treated as an agent that interacts with various users and receives user feedback over time. However, an agent considering instant user feedback of a single request may allocate live streams greedily in a short video feed and eventually affect user retention. Therefore, the optimal live stream injection control problem is modeled as an infinite request level CMDP to maximize the cumulative live stream reward and satisfy the platform-level constraint, for example, live stream allocation cannot reduce the total app usage duration.
Formally, we define live stream injection CMDP as a tuple of six elements $\left(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\mathcal{C}\right)$:
@@
  • Constraint $\mathcal{C}$: The platform-level constraint is that the live stream watch time should be longer than the average video watch time per request as much as possible. This prevents a negative impact on app usage duration, which could happen if the live stream watch time is shorter. By meeting this constraint, we ensure that live streams boost users’ long-term app engagement instead of reducing their time spent on it.
@@
Overall, the optimization objective of aforementioned optimal live stream injection control problem is shown as follows:
@@
$J_{R}(\pi)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$
$J_{C}(\pi)=\frac{1}{B}y_{v}-y_{l}$
where $\pi(a_{t}|s_{t})$ is the live stream injection policy to be optimized, $y_{l}$ is the live stream watch time, $y_{v}$ is the total video watch time, and $B$ is the number of videos in a request.
@@
Inspired by the RCPO (Tessler et al., 2018) algorithm, the primal-dual optimization problem can be transformed into an equivalent unconstrained problem with penalized reward function:
@@
In practice, we solve the unconstrained MDP problem and the reward function is simplified as follows:
(4) $r(s_{t},a_{t})=y_{l}-\frac{\lambda}{B}y_{v}$
where $\lambda$ is a hyper-parameter to carefully control the final live stream injection ratio. From the perspective of constrained reinforcement learning, the term $\frac{\lambda}{B}y_{v}$ can be viewed as an implicit cost constraint.
@@ 3.2. Multi-Group State Decomposition Module
Since live stream content is not equally appealing to all users of the short video app Kwai, it is crucial to inject live streams selectively into the short video feed. Otherwise, excessive live stream exposures may disrupt a user’s interest in short videos, leading to a decrease in overall app usage duration. In practice, user interaction data for live streams are sparser and noisier compared to that for short videos, making it challenging for RL algorithms to learn an optimal live stream injection policy for each individual user. Directly learning a policy from sparse and noisy user feedback is often infeasible, since it will inadvertently inject excessive live streams in the short video feed and highly affected the user engagement.
@@ 4.1. Dataset
Due to the lack of publicly available datasets for recommendation decisions involving live streaming injections in short video feeds, we collect an offline dataset from our recommendation system with 100,000 users through random sampling over a full day for offline evaluation, ensuring that the proportions of different user groups are similar to those in real online data. In total, we have over 1,800,000 samples, which are divided into training and test sets with a 4:1 ratio. In addition, the online environment contains more than 100 million app users for AB-Tests.
@@ 4.2.1. Baselines
Note that in our optimal live stream injection control problem, the action is discrete. Therefore, we introduce the Straight-Through Gumbel Softmax (Jang et al., 2016) technique in TD3 and TD3-BC.
@@ 4.7. Online A/B Test Experiment
We compare the improvements of different methods in terms of live stream daily active users (DAU), live stream watch time, video watch time, app usage duration and user retention relative to the baseline. The baseline uses the SAC framework for live stream injection. The results of the online A / B test are shown in Table 3.
Note that the Dummy method injects a live stream for each request. We can see that SL-MGAC achieves the largest improvement in live stream watch time while also increasing the app usage duration and user retention than other methods. It demonstrates that SL-MGAC is more effective in maximizing long-term rewards when injecting live streams in response to user requests. Although SL-MGAC-0 shows a greater improvement in live stream DAU compared to SL-MGAC, it tends to be more greedy in injecting live streams, which may negatively impact the long-term user experience for most users. Moreover, the dummy method can greatly improve live streaming DAU and watch time, but significantly affect app usage time and user retention.
@@ 4.8. Online Model Stability
We evaluate the online model stability between the SAC-based baseline and SL-MGAC. As shown in Fig. 6, the live stream injection ratio fluctuates drastically throughout the day for both models. However, SL-MGAC exhibits greater stability than the baseline, with a smaller amplitude. Note that the live stream injection ratio refers to the proportion of requests with action $a=1$.
Figure 6. The trend curve of online live stream injection ratio during a whole day.
We calculate the amplitudes of the live stream injection ratio within sliding time windows of 20 minutes. The amplitude density is shown in Fig. 8 in Appendix B. The results indicate that SL-MGAC has a smaller mean amplitude and fewer outliers. In contrast, the baseline model exhibits larger amplitude outliers, as highlighted in the red dashed box, demonstrating the superior online model stability of SL-MGAC.
5. Discussion
The intrinsic instability and divergence of RL limits its application in real-world industrial recommendation systems. The proposed SL-MGAC method improves the robustness and stability of the RL model and achieves significant offline and online performance in the live stream and short video mixed recommendation scenario. However, SL-MGAC only learns from data from adjacent states $(s_{t},a_{t},r_{t},s_{t+1},a_{t+1},r_{t+1})$, which limits its ability to learn information in much longer sequences, such as the daily-level user preference, rather than session-level user preference, since the capture of daily user preference information is very helpful for improving user engagement and user retention. Learning live stream policy from a long sequence of user behaviors would lead to our future work.
Furthermore, we emphasize that more sophisticated methods can also be employed to divide users into multiple groups and provide prior information about user preference, such as traditional clustering methods, e.g. k-means, neural clustering of VQ-VAE (Van Den Oord et al., 2017), online clustering of SwAV (Caron et al., 2020) by the Sinkhorn-Knopp algorithm (Cuturi, 2013), etc. We find that the simple MG-SD module achieves significant performance and fast real-time response spped in practice; we will lead these complicated group partitioning methods in our further research.
@@ 7. Conclusion
In this work, we propose a novel Supervised Learning-enhanced Multi-Group Actor-Critic algorithm (SL-MGAC) to optimize request-level live stream injection policies in the challenging context of a live stream & short video mixed recommendation system. Unlike existing RL methods for recommendation systems, our approach focuses on enhancing the stability and robustness of the RL model. Specifically, we introduce a multi-group state decomposition module to reduce prediction variance and seamlessly integrate multi-task reward learning with traditional critic learning to constrain Q-value estimation. In practice, we aim to minimize the risk of policy deterioration or model collapse, thereby enabling the proposed SL-MGAC method to be successfully deployed in large-scale industrial recommendation systems.

### v4

@@ Abstract.
In the context of a short video & live stream mixed recommendation scenario, the live stream recommendation system (RS) decides whether to allocate at most one live stream into the video feed for each user request. To maximize long-term user engagement, it is crucial to determine an optimal live stream policy for accurate live stream allocation. The inappropriate live stream allocation policy can significantly affect the duration of the usage app and user retention, which ignores the long-term negative impact of live stream allocation. Recently, reinforcement learning (RL) has been widely applied in recommendation systems to capture long-term user engagement. However, traditional RL algorithms often face divergence and instability problems, which restricts the application and deployment in the large-scale industrial recommendation systems, especially in the aforementioned challenging scenario. To address these challenges, we propose a novel Supervised Learning-enhanced Multi-Group Actor Critic algorithm (SL-MGAC). Specifically, we introduce a supervised learning-enhanced actor-critic framework that incorporates variance reduction techniques, where multi-task reward learning helps restrict bootstrapping error accumulation during critic learning. Additionally, we design a multi-group state decomposition module for both actor and critic networks to reduce prediction variance and improve model stability. We also propose a novel reward function to prevent overly greedy live stream allocation. Empirically, we evaluate the SL-MGAC algorithm using offline policy evaluation (OPE) and online A/B testing. Experimental results demonstrate that the proposed method not only outperforms baseline methods under the platform-level constraints but also exhibits enhanced stability in online recommendation scenarios.
@@ 1. Introduction
We consider a challenging sequential decision making scenario in a short video & live stream blended recommendation system, as shown in Fig. 1. The system consists of three components: a live streaming recommendation system, a short video recommendation system, and a blending server. For each user request with timestamp $t$, the live stream recommendation system decides whether to inject the recommended live stream into the video feed, while the short video recommendation system suggests $B$ (with $B<10$) videos. The blending server then mixes and rearranges the single live stream with the $B$ videos to form the final recommendation list. Our objective is to find a personalized optimal live stream allocation policy that not only maximizes user long-term engagement with live streams, but also meets the constraint of not causing a negative impact on the app usage duration and user retention, which can be modeled as an infinite request-level Constrained Markov Decision Process (CMDP) (Altman, 2021).
Currently, deep reinforcement learning has shown great potential in learning policies to maximize long-term rewards in various research domains, such as computer vision, robotics, natural language processing, gaming, and recommendation systems (Afsar et al., 2022; Li, 2017). In the context of recommendation systems, reinforcement learning is applied to optimize long-term user engagement (Zou et al., 2019) and improve user retention (Cai et al., 2023). Numerous RL applications have been proposed for real-world recommendation systems, including slate recommendation (Ie et al., 2019; Deffayet et al., 2023; Liu et al., 2023), personalized search ranking (Miao et al., 2021), and advertisement allocation (Liao et al., 2022).
However, the intrinsic issues of divergence and instability associated with traditional reinforcement learning algorithms (RL) (François-Lavet et al., 2015; Fujimoto et al., 2018; Kumar et al., 2019; Dasagi et al., 2019) are significantly exacerbated in the short video & live stream mixed recommendation system. A primary possible reason for this is the drastic fluctuations in both the live stream supply scale and user interaction behaviors over time, as shown in Fig. 7 in Appendix B. RL models often struggle to learn effective policies from data that exhibit such high variance across both time scales and user scales. In practice, we observe that RL models frequently encounter issues of policy deterioration or model collapse (Dohare et al., 2023).
@@
The main contributions of this paper are as follows.
  • We introduce SL-MGAC, a novel RL model that outperforms existing methods in maximizing the user engagement of live stream while satisfying the platform-level constraints on app usage duration and user retention.
@@ 2. Problem Formulation
As shown in Fig. 1, in the short video & live stream mixed recommendation system, the live stream recommendation system is treated as an agent that interacts with various users and receives user feedback over time. However, an agent considering instant user feedback of a single request may allocate live streams greedily in a short video feed and eventually affect user retention. Therefore, the optimal live stream allocation control problem is modeled as an infinite request level CMDP to maximize the cumulative live stream reward and satisfy the platform-level constraint, for example, live stream allocation cannot reduce the total app usage duration.
Formally, we define live stream allocation CMDP as a tuple of six elements $\left(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma,\mathcal{C}\right)$:
@@
  • Constraint $\mathcal{C}$: The platform-level constraint is an instantaneous constraint, that is, the live stream watch time should be longer than the average video watch time per request as much as possible. This prevents a negative impact on app usage duration, which could happen if the live stream watch time is shorter. By meeting this constraint, we ensure that live streams boost users’ long-term app engagement instead of reducing their time spent on it.
@@
Overall, the concrete reward function $r(s_{t},a_{t})$ and constraint function $c(s_{t},a_{t})$ is shown below:
(1) $r(s_{t},a_{t})$ $=y_{l}$
$c(s_{t},a_{t})$ $=\frac{1}{B}y_{v}-y_{l}$
where $y_{l}$ is the live stream watch time, $y_{v}$ is the total video watch time, and $B$ is the number of videos in a request.
The optimization objective of aforementioned optimal live stream allocation control problem is shown as follows:
@@
where $J_{R}(\pi)=\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$, $J_{C}(\pi)=c(s_{t},a_{t}),\forall t\in[0,\infty)$, $\pi(s_{t},a_{t})$ is the live stream allocation policy to be optimized.
@@
Inspired by the RCPO (Tessler et al., 2018) algorithm, the primal-dual optimization problem can be transformed into an equivalent unconstrained problem with penalized reward function by substituting into $r(s_{t},a_{t})$ and $c(s_{t},a_{t})$ of Eq. 1:
@@
In practice, we solve the unconstrained MDP problem with a simplified reward function as follows:
(5) $\max_{\pi}\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}\left(y_{l}-\frac{\lambda}{B}y_{v}\right)\right]$
where $\lambda$ is a hyper-parameter to carefully control the final live stream allocation ratio. From the perspective of constrained reinforcement learning, the term $\frac{\lambda}{B}y_{v}$ can be viewed as an implicit cost constraint. Note that we leave the further discussion of the rationality of the above reward function in the Appendix C.
@@ 3.2. Multi-Group State Decomposition Module
Since live stream content is not equally appealing to all users of the short video app Kwai, it is crucial to inject live streams selectively into the short video feed. Otherwise, excessive live stream exposures may disrupt a user’s interest in short videos, leading to a decrease in overall app usage duration. In practice, user interaction data for live streams are sparser and noisier compared to that for short videos, making it challenging for RL algorithms to learn an optimal live stream allocation policy for each individual user. Directly learning a policy from sparse and noisy user feedback is often infeasible, since it will inadvertently inject excessive live streams in the short video feed and highly affected the user engagement.
@@ 4.1. Dataset
Due to the lack of publicly available datasets for recommendation decisions involving live streaming allocations in short video feeds, we collect an offline dataset from our recommendation system with 100,000 users through random sampling over a full day for offline evaluation, ensuring that the proportions of different user groups are similar to those in real online data. In total, we have over 1,800,000 samples, which are divided into training and test sets with a 4:1 ratio. In addition, the online environment contains more than 100 million app users for AB-Tests.
@@ 4.2.1. Baselines
Note that in our optimal live stream allocation control problem, the action is discrete. Therefore, we introduce the Straight-Through Gumbel Softmax (Jang et al., 2016) technique in TD3 and TD3-BC.
@@ 4.7. Online A/B Test Experiment
We compare the improvements of different methods in terms of daily active live stream users (DAU), live stream watch time, video watch time, app usage duration, and user retention relative to the baseline. The baseline uses the SAC framework for live stream allocation. The results of the online A / B test are shown in Table 3.
Note that the Dummy method injects a live stream for each request. We can see that SL-MGAC achieves a significant improvement in live stream watch time while also increasing the app usage duration and user retention than other methods. This shows that SL-MGAC is more effective in maximizing long-term live stream rewards while avoiding the negative impact on app usage duration and user retention. Although SL-MGAC-0 shows a greater improvement in live stream DAU compared to SL-MGAC, it tends to be more greedy in injecting live streams, which may affect the long-term user experience for most users. Moreover, although the dummy method can greatly improve DAU and live stream viewing time, it significantly affects app usage time and user retention.
@@ 4.8. Online Model Stability
We evaluate the online model stability between the SAC-based baseline and SL-MGAC. As shown in Fig. 6, the live stream allocation ratio fluctuates drastically throughout the day for both models. However, SL-MGAC exhibits greater stability than the baseline, with a smaller amplitude. Note that the live stream allocation ratio refers to the proportion of requests with action $a=1$.
Figure 6. The trend curve of online live stream allocation ratio during a whole day.
We calculate the amplitudes of the live stream allocation ratio within sliding time windows of 20 minutes. The amplitude density is shown in Fig. 8 in Appendix B. The results indicate that SL-MGAC has a smaller mean amplitude and fewer outliers. In contrast, the baseline model exhibits larger amplitude outliers, as highlighted in the red dashed box, demonstrating the superior online model stability of SL-MGAC.
@@ 6. Conclusion
In this work, we propose a novel Supervised Learning-enhanced Multi-Group Actor-Critic algorithm (SL-MGAC) to optimize request-level live stream allocation policies under the platform-level constraints on app usage duration and user retention in the challenging context of a live stream & short video mixed recommendation system. Specifically, we introduce a multi-group state decomposition module to reduce prediction variance and seamlessly integrate multi-task reward learning with traditional critic learning to constrain Q-value estimation. Compared with existing RL methods for recommendation systems, our approach not only focuses on improving online performance, but also improves the stability and robustness of the RL model. In practice, we minimize the risk of policy deterioration or model collapse, thereby enabling the proposed SL-MGAC method to be successfully implemented in large-scale industrial recommendation systems. Currently, we are trying to apply the SL-MGAC methods to other similar recommendation scenarios, including E-commerce and advertising. We also aim to continuously improve the performance of SL-MGAC in further research. Learning live stream policy from a long sequence of user behaviors $(s_{t},a_{t},r_{t})$ and exploring more refined and complex user state representations would lead to our future work.
7. Acknowledgment
The authors acknowledge Bai Wei and Chen Xiaoshuang for proposing detailed modification advice to help us improve the quality of this manuscript.